Shadows of diagrams and Euler-Zagier sums

GoF

a Institut Galilée, LIPN, CNRS UMR 7030 99 Av. J.-B. Clement, F-93430 Villetaneuse, France

b H. Niewodniczański Institute of Nuclear Physics, Polish Academy of Sciences ul. Eliasza-Radzikowskiego 152, PL 31342 Kraków, Poland

c Laboratoire de Physique Théorique de la Matière Condensée Université Pierre et Marie Curie, CNRS UMR 7600 Tour 24 - 2ième ét., 4 pl. Jussieu, F 75252 Paris Cedex 05, France

e The Open University, Physics and Astronomy Department Milton Keynes MK7 6AA, United Kingdom

E-mail:

ghed@lipn-univ.paris13.fr, a.i.solomon@open.ac.uk
pawel.blasiak@ifj.edu.pl, andrzej.horzela@ifj.edu.pl,
penson@lptl.jussieu.fr,

Abstract. We

Contents

1 Introduction 2

2 The deformed algebra of labelled diagrams 2 2.1 How these diagrams arise . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 2.2 Codings and the deformed algebra . . . . . . . . . . . . . . . . . . . . . . 2 2.3 LBELL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

3 Euler-Zagier sums 2 3.1 Definitions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 3.2 Shadows and interpretations . . . . . . . . . . . . . . . . . . . . . . . . . 2 3.3 General picture . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

4 Conclusion 2 CONTENTS 2

1. Introduction

2. The deformed algebra of labelled diagrams

2.1. How these diagrams arise

2.2. Codings and the deformed algebra

2.3. LBELL

3. Euler-Zagier sums

3.1. Definitions

3.2. Shadows and interpretations

3.3. General picture

4. Conclusion

Discuss the self-dual deformation.

References

[1] C. M. Bender, D. C. Brody, and B. K. Meister, Quantum field theory of partitions, J. Math. Phys. Vol 40 (1999)

[2] Bourbaki N., Algèbre, chapitre III, Hermann ??

[3] Cartier P., Séminaire “Sopus Lie”, 2ème année, Faculté des Sciences de Paris (1955-56)

[4] V. Chari, A. Pressley, A guide to quantum groups. Cambridge (1994).

[5] L. Comtet, Analyse Combinatoire, PUF, Tome 1 (1970)

[6] L. Comtet, Analyse Combinatoire, PUF, Tome 2 (1970)

[7] P. Flajolet, http://algo.inria.fr/flajolet/

[8] L. Foissy, Isomorphisme entre l’algèbre des fonctions quasi-symétriques libres et une algèbre de Hopf des arbres enracinés décorés plans, personal communication.

[9] L. Foissy, Les algèbres de Hopf des arbres enracinés decorés, PhD Memoir, Reims University (2002).

[10] A.I. Solomon, P. Blasiak, G. Duchamp, A. Horzela and K.A. Penson, Combinatorial Physics, Normal Order and Model Feynman Graphs, Proceedings of the Symposium ’Symmetries in Science XIII’, Bregenz, Austria, 2003, B. Gruber, G. Marmo and N. Yoshinaga (eds.), p.527 (Kluwer Academic Publishers 2004)

arXiv : quant-ph/0310174

[11] A.I. Solomon, G. Duchamp, P. Blasiak, A. Horzela and K.A. Penson, Normal Order: Combinatorial Graphs Quantum Theory and Symmetries, Proceedings of the 3rd International Symposium P.C. Argyres, T.J. Hodges, F. Mansouri, J.J. Scanio, P. Suranyi, and L.C.R. Wijewardhana (eds.), p.398 (World Scientific Publishing 2004)

arXiv:quant-ph/0402082

[12] G. H. E. Duchamp, P. Blasiak, A. Horzela, K. A. Penson, A. I. Solomon, Feynman graphs and related Hopf algebras, Journal of Physics: Conference Series, SSPCM’05, Myczkowce, Poland. arXiv : cs.SC/0510041

[13] G. Duchamp, A.I. Solomon, K.A. Penson, A. Horzela and P. B lasiak, One-parameter groups and combinatorial physics, Proceedings of the Symposium Third International

CONTENTS 3

Workshop on Contemporary Problems in Mathematical Physics (COPROMAPH3) (Porto-Novo, Benin, Nov. 2003), J. Govaerts, M. N. Hounkonnou and A. Z. Msezane (eds.), p.436 (World Scientific Publishing 2004)

arXiv: quant-ph/04011262

[14] A.I. Solomon, G. Duchamp, P. Blasiak, A. Horzela and K. A. Penson, Partition functions and graphs: A combinatorial approach, Proceedings of the XI International Conference on Symmetry Methods in Physics (SYMPHYS-11) (Prague, Czech Republic, June 2004), C. Burdik, O. Navratil, and S. Posta (eds.) (JINR Publishers, Dubna, 2004)

arXiv:quant-ph/0409082

[15] A. Horzela, P. Blasiak, G. Duchamp, K. A. Penson and A.I. Solomon, A product formula and combinatorial field theory, Proceedings of the XI International Conference on Symmetry Methods in Physics (SYMPHYS-11) (Prague, Czech Republic, June 2004), C. Burdik, O. Navratil, and S. Posta (eds.) (JINR Publishers, Dubna)

arXiv:quant-ph/0409152

[16] P. Blasiak, A. Horzela, K. A. Penson, G. H. E. Duchamp, A.I. Solomon, Boson normal ordering via substitutions and Sheffer-Type Polynomials, Phys. Lett. A 338 (2005) 108

[17] P. Blasiak, K. A. Penson, A.I. Solomon, A. Horzela, G. H. E. Duchamp, Some useful formula for bosonic operators, Jour. Math. Phys. 46 052110 (2005).

[18] G. Duchamp, F. Hivert, J. Y. Thibon, Non commutative functions VI: Free quasisymmetric functions and related algebras, International Journal of Algebra and Computation Vol 12, No 5 (2002).

[19] S. A. Joni, G.-C.Rota, Colgebras and Bialgebras in Combinatorics, Studies in Applied Mathematics 61, 93-139 (1979).

[20] D. Knuth, The art of computer programming Tome I. Addison-Wesley (1981)

[21] D. Kreimer, Knots nad Feynman diagrams, Cambridge.

[22] J. Milnor, J. Moore, On the structure of Hopf Algebras, Ann. of Math. 81 (1965), 211-264.

[23] R. P. Stanley, Enumerative Combinatorics, vol. 2, Cambridge Univ. Press (2004).

[24] S. Weinzierl, Hopf algebra structures in particle physics, Eur. Phys. J. C 33 (2004) arXiv:

hep-th/0310124